# Searching from Area to Point: A Semantic Guided Framework with Geometric Consistency for Accurate Feature Matching

### v5

@@ Abstract
Feature matching is a crucial technique in computer vision. A unified perspective for this task involves treating it as a search problem, with the objective of developing an efficient search strategy to narrow down the search space to achieve point matches between images. The search space is a key element which affects the search strategy design, but it is often not well defined in current approaches, leading to limited accuracy in matching. This paper, thus, proposes a novel search space for precise feature matching, named semantic area matches, which consists of matched image areas containing prominent semantic. This search space facilitates point matching by reducing the matching difficulty and enables recent detector-free matching methods to handle high-resolution input. For the search strategy corresponding to this search space, we propose a hierarchical feature matching framework called Area to Point Matching (A2PM), which involves identifying semantic area matches between images and subsequently performing point matching on these area matches. Additionally, we propose the Semantic and Geometry Area Matching (SGAM) method to implement this framework, which leverages semantic priors and geometric consistency to establish precise area and point matches between images. By integrating the SGAM method with off-the-shelf detector-free matchers to adopt the A2PM framework, our proposed method demonstrates significant performance improvement in extensive point matching (up to $+29.16\%$) and pose estimation (up to $+11.04\%$) experiments. The code is publicly available at https://github.com/Easonyesheng/SGAM.
@@
Feature matching, pose estimation, Correspondence Estimation, Epipolar Geometry
@@ I Introduction
Feature matching is a fundamental task in computer vision, which serves as the basis of a wide range of vision applications, such as simultaneous localization and mapping [1], structure from motion [2] and image alignment [3]. Despite its status as a well-studied task, accurately determining the projections of a single 3D point in two different viewpoints continues to pose challenges. These challenges arise from potential extreme viewpoints, light variations, repetitive patterns, and motion blur, all of which contribute to the limited accuracy in the downstream two-view geometry estimation.
In pursuit of high precision feature matching, this paper proposes a unified searching perspective for this task. Given the source point in one image, determining the correspondence in the other image involves searching for the target point from a specific set (the search space) by a carefully designed approach (the search strategy). Therefore, he search space and search strategy are crucial for this searching task. Specifically, the search space is a set containing the targets, which vary through different search phases. It functions as the input or output of the search strategy, which aims to narrow down the input search space to the target one. Hence, a well-defined search space accompanied by an effective search strategy ensure high precision in searching (Fig. 1).
Current feature matching methods are divided into detector-based and detector-free frameworks depending on whether a detector is used. However, taking this unified searching perspective into consideration, the fundamental disparity primarily arises from the distinct search spaces, which leads to the employment of different search strategies. In the detector-based framework [4, 5], the search space is first reduced from the entire image to the keypoint set, from which the correspondences are subsequently established. On the other hand, patch-level correspondences are utilized as the mid-level search space in the detector-free framework [6, 7]. These correspondences are obtained through dense feature comparison between images. Then, point matches are refined from these matched patches.
Figure 1: Performance comparison on MatterPort3D. The dashed boxes are area matches obtained by SGAM, providing current point matchers with reasonable search space (left, where the original LoFTR failed) and high input image resolution (right) to achieve significant performance improvement.
While finding point matches from keypoint set is easy in the detector-based framework, detecting keypoints even with deep CNN [8, 9] in the image, however, suffers from inaccurate and failed detection caused by rapid change in viewpoint, varied illumination, complex 3D geometry and motion blur. Thus the keypoint set is not a suitable mid-level search space as it is difficult to obtain reliably. The detector-free framework [10] adopts the dense search strategy in the entire image and takes patch-level matches as the mid-level search space, where the point-wise match refinement is performed. Attention mechanism is utilized in this effective coarse-to-fine framework to endow perception of global context and achieves SOTA performance [11, 7, 12, 13]. Nevertheless, the patch matches lack clear definitions related to image contents, resulting in expensive feature computations to obtain them. Therefore, in the coarse level, the full-image feature extraction and attention computation have to be implemented in size-reduced images for acceptable computation cost, in which much detailed information is lost, limiting the matching accuracy. Recently, another two-stage matching method, MKPC [14], achieves the co-visible area between images as the mid-level search spaces by off-the-shelf point matcher, but the individual co-visible area is too rough and its accuracy suffers from stage-one matching. PATS [15] utilizes evenly cropped image patches as the mid-level search space, but achieving these matches necessitates multiple iterations, thereby incurring additional computational costs. In summary, existing feature matching methods suffer from a lack of appropriate search space and corresponding search strategy, limiting the matching precision.
@@
To tackle this problem, we propose to utilize semantic information in images to establish a novel search space for feature matching. Firstly, the study of semantic understanding in images is well-established, and recent semantic segmentation methods [16] demonstrate exceptional accuracy and generalization, naturally generating multiple areas containing substantial semantic information inside the image, termed as semantic areas. Due to the invariance of semantic to viewpoint, illumination, and scale [17, 18, 19], precise semantic segmentation enables seamless matching of semantic areas between images. Thus, the semantic area matches are easily accessible. On the other hand, the semantic areas facilitate inside point match searching, since distinct image features that aid in accurate point matching are typically concentrated within these areas. Meanwhile, the area-based computation is more feasible compared to entire-image-based computation. Hence, utilizing this semantic area matches as the initial search space for detector-free methods leads to more reliable point matching and allows for high-resolution inputs, ultimately improving the matching performance. Therefore, the semantic area matches are proper to serve as both the mid-level search space for feature matching and the primary search space for off-the-shelf detector-free methods.
In order to align the proposed search space with an efficient search strategy, we propose a hierarchical feature matching framework: Area to Point Matching (A2PM, Fig.2 top). This framework contains two search phases: matching semantic areas from the entire images and matching points from the matched areas. For the first phase, we introduce Semantic Area Matching (SAM), utilizing the semantic prior to efficiently detect and match semantic areas between images. In specific, SAM leverages semantic segmentation images as input to match two types of image areas: semantic object areas and semantic intersection areas. Through a process of detection, description, and matching, putative area matches are achieved. However, the use of semantic abstraction in SAM, which omits local details, can introduce semantic ambiguity during matching, particularly when distinct instances coexist within the image. Therefore, SAM also identifies doubtful areas that cannot be confidently matched.
@@
Besides doubtful areas, the scene-dependent semantic ambiguity may also lead to erroneous area matches in SAM, negatively impacting feature matching. Therefore, we turn to the inherent geometry property of area matches. Specifically, considering that area matches in images signifies the same 3D entity, the constraint of epipolar geometry naturally applies, regardless of the scene. Furthermore, area matches in the same image pair must adhere to the same constraint, named geometric consistency, which can be employed to address the semantic ambiguity. Thus, we establish the geometric consistency of area matches by utilizing fundamental matrices of point matches within these areas. It enables the proposed Geometry Area Matching (GAM) to integrate an off-the-shelf point matcher for accurate determination of area matches. In practice, the GAM first predicts true area matches from doubtful candidates generated by SAM (GAM Predictor, GP). Subsequently, all area matches undergo filtering by the GAM Rejector (GR) to identify matches with superior geometric consistency. The point matches within accurate area matches are obtained concurrently, thereby completing the second search phase of the A2PM. Furthermore, to handle the correspondence aggregation issue in less-semantic scenes, the Global Match Collection module (GMC) is incorporated in GAM, which involves collecting additional point matches globally based on the geometry consistency of inside-area matches. The GMC ensures the generation of well-distributed matches, advantageous for downstream tasks. Through the combination of SAM and GAM, our Semantic and Geometry Area Matching (SGAM, Fig. 2 bottom) is capable of achieving accurate area and point matching between images.
In sum, this work presents the following contributions:
@@
    A well-defined search space for feature matching, i.e. the semantic area matches, is proposed under a hierarchical feature matching framework: A2PM, which narrows down the matching search space to area matches with prominent semantic for precise point matching.
@@
    Our method provides detector-free methods with a suitable search space and high-resolution input (Fig. 1), thereby achieving notable improvements in matching precision (up to $+29.16\%$) and pose estimation performance (up to $+11.04\%$), with semantic segmentation achieved by SEEM as the input.
TABLE I: Symbol Table. The table provides a comprehensive list of symbols used in the paper, and brief descriptions for each.
Symbol Description
$I_{i}$ Input image pair, $i\in\{0,1\}$
[…]
indicating the $l$-th area matching possibility.
[…]
=$=\{(q_{i}^{m},p_{i}^{m})\}_{m}^{M}$ $i$ is the index of point match set,
$m$ is the index of point pair inside the set.
[…]
calculated by $F_{i}$ and a point pair $(q_{i}^{m},p_{i}^{m})$.
[…]
combined with point matcher $PM$.
[…]
combined with point matcher $PM$.
[…]
with index $l$ related to $\pi_{l}(i)$.
[…]
in the image pair.
[…]
@@ II Related Work
II-A Detector-based Matching
Detector-based matching relies on detected keypoints and their descriptors [4, 20]. Through the nearest neighbor search based on descriptor distances, point matches can be established between images. In the age of deep learning, recent work [9, 5, 21, 22] utilize deep CNN to achieve better learning feature. Specifically, SuperPoint [5] is early in providing feature detection and description networks and outperforms conventional methods. Subsequent work [23, 9, 8] leverage a unified network to detect and describe feature. At the same time, detached learning detection [24, 25, 26] and description [27, 28, 29] are proposed as well. After feature detection, point match searching and outlier rejection are also advanced by recent learning methods [30, 31, 32, 33, 34, 35]. Essentially, this framework relies on extremely fine-grained search space establishment, i.e. the feature points, to achieve accurate point matching. However, feature point detection poses significant challenges in scenes with low texture, repetitive patterns, extreme changes in illumination and scale, which ultimately leads to a decline in performance. In contrast, our A2PM framework appropriately reduces the search space to semantic area matches, which are more robust due to their semantic invariance.
II-B Detector-free Matching
In order to avoid detection failure, detector-free framework is proposed [36, 10, 37], aiming at jointly trainable feature detection, matching and outlier rejection to establish point matches directly from image pairs. Initially, 4D CNN is used to extract dense image features [36, 38, 39]. Recent DKM [40] constructs a Gaussian Process using CNN and achieves leading performance. Meanwhile, owing to limited receptive field [10] of CNN, some methods [10, 37] incorporate Transformer [41] to process dense feature extracted by CNN, attaining SOTA performance. However, the global search space in this framework leads to redundant computation, limiting the resolution of input images. Furthermore, the redundant search space introduces noise from non-overlapping areas in the image pair, resulting in degraded precision of attention operations. As a solution, we propose the A2PM framework, which performs area matching in images prior to point matching. Due to the higher resolution and less noise in area matches, subsequent inside-area point matching precision is improved.
@@ II-C Coarse-to-Fine Matching
To avoid high computation costs due to the redundant full-image search space in detector-free matching, coarse-to-fine methods are proposed. In LoFTR [10], patch-level correspondences are established through feature attention of the coarse level matching, which serves as the initial search space for the fine level matching. This approach significantly reduces time consumption compared to the vanilla Transformer [37]. Subsequent works [11, 7, 12, 13] have improved the attention mechanism to construct a more efficient and precise coarse search space. However, this patch-level search space lacks a clear association with image apparent features or high-level semantics, thus requiring expensive feature comparison for correspondence establishment. The accuracy issue in the coarse level also persists due to low input resolution. PATS [15] is proposed to extract more accurate features from equally cropped image patches with high resolution, but finding patch matches requires multiple iterations. Recent MKPC [14] provides another two-stage feature matching strategy, which utilizes stage-one point matching on the entire images to achieve the co-visible area. Then the stage-two point matching is performed inside it. However, individual co-visible area matching is too rough for precise point matching and its accuracy is entangled with the stage-one point matching. In contrast, our method offers a well-defined search space, i.e. the semantic area matches, along with an independent area matching method utilizing both semantic and geometry priors. It is able to provide a fine-grained search space for detector-free matchers, improving their matching performance.
III Formulation
In this section, we define the proposed A2PM framework and establish a foundation of geometry consistency for our method. Symbols used in this section can be found in Tab. I.
III-A Framework Formulation
@@
Generally, given an image pair ($I_{0},I_{1}$), an Area Matching method $AM$ and a detector-free Point Matching method $PM$, our A2PM framework ($\mathcal{M}_{A}$) is formulated as:
$\mathcal{P}=\mathcal{M}_{A}(I_{0},I_{1},AM,PM)$ (1)
The output $\mathcal{P}=\{q^{m},p^{m}\}_{m}^{M}$ is the set of point matches.
@@
Specifically, we formulate the first search phase from the image pair to area matches as follows. Suppose areas in $I_{0},I_{1}$ are respectively $\{\alpha_{i}\}_{i}$ and $\{\beta_{j}\}_{j}$. The area match is represented as $\mathcal{A}_{i,j}=(\alpha_{i},\beta_{j})$. Therefore, $N$ pairs of area matches can be achieved by our $AM$:
$\{\mathcal{A}_{i,\pi(i)}\}_{i}^{N}=AM(I_{0},I_{1})$ (2)
@@
where $\pi(i):\mathbb{R}\rightarrow\mathbb{R}$ is the index mapping between matched areas. Then, point matches inside area matches can be found:
$\mathcal{P}=PM(\{\mathcal{A}_{i,\pi(i)}\}_{i}^{N})$ (3)
@@
Due to the proper search space of semantic area matches, the matching accuracy of $PM$ is improved. Moreover, the area matches possess higher resolution than the original input of $PM$, leading to more precise matches.
Figure 3: Semantic Area Matching (SAM). SAM divides image areas into two types to process. For both types of areas, area detection and description are performed first, and then area matching is based on specially-designed descriptor distance minimization.
III-B Geometry Consistency Formulation

### v6

@@ Abstract
Feature matching plays a pivotal role in computer vision applications. To achieve efficient and accurate matching, current methods commonly employ a coarse-to-fine strategy, which establishes an intermediate search space preceding point matches. However, the difficulty in establishing dependable intermediate search spaces poses a limitation on the overall matching performance of existing feature matching methods. To address this issue, this paper proposes the integration of robust semantic priors in the intermediate search space and introduces a semantic-friendly search space called semantic area matches for precise feature matching. The semantic area matches comprise matched image areas with significant semantic content, which can be robustly attained due to the semantic invariance against matching noise. Moreover, it facilitates point matching by reducing the redundancy and enables high-resolution input. To adopt this search space, we introduce a hierarchical feature matching framework called Area to Point Matching (A2PM), which involves identifying semantic area matches between images and subsequently conducting point matching on these area matches. Furthermore, we present the Semantic and Geometry Area Matching (SGAM) method to implement this framework, which leverages semantic priors and geometric consistency to establish precise area and point matches between images. Through the adoption of the A2PM framework, SGAM demonstrates substantial and consistent enhancements in the performance of sparse, semi-dense, and dense point matchers in extensive point matching (up to $+29.16\%$) and pose estimation (up to $+13.01\%$) experiments. The code is publicly available at https://github.com/Easonyesheng/SGAM.
@@
Feature matching, pose estimation, Epipolar Geometry
@@ I Introduction
Feature matching is a fundamental task in computer vision, which serves as the basis of a wide range of vision applications, such as simultaneous localization and mapping [1], structure from motion [2] and image alignment [3]. Despite its status as a well-studied task, accurately determining the projections of a single 3D point in two different viewpoints continues to pose challenges. These challenges arise from matching noises, such as potential extreme viewpoints, light variations, repetitive patterns, and motion blur, all of which result in the limited matching accuracy.
Figure 1: The proposed semantic-friendly search space of feature matching. This search space, termed as semantic area matches, can be robustly established by leveraging semantic invariance, thereby reducing redundant computations in feature matching. As a result, fine-grained search spaces can be reliably established within the area image pairs, consequently enhancing matching accuracy.
Current feature matching methods are divided into sparse, semi-dense and dense methods [4]. Despite differences in specific techniques, narrowing the search space by means of hierarchical matching is the consensus of these methods. Typically, these methods begin by establishing intermediate search spaces for point matching between images, followed by the point matching within these spaces. Specifically, in sparse methods [5, 6], the keypoint set is first detected in images, from which correspondences are subsequently achieved. While finding point matches from the keypoint set is easy in the sparse methods, detecting keypoints even with deep CNN [7, 8], however, suffers from inaccurate and failed detection caused by matching noises. In semi-dense methods [9, 10], sparse image patch matches are initially found by dense feature comparison, with point matches refined from these patches. Similarly, dense methods [11, 12] progressively refine dense warps from coarse to fine feature maps, utilizing dense patch matches between images as the intermediate search space. However, the patch matching relies on dense feature comparison, which leads to error-prone computation on irrelevant features and limited input resolution, consequently impacting the overall accuracy of semi-dense and dense methods.
To address the redundant computation, two-stage matching methods [13, 14] propose to achieve the co-visible area (or region) match between images as an intermediate search space of feature matching. Nonetheless, establishing this search space still heavily relies on feature comparison and suffers from matching noises. On the other hand, PATS [15] proposes a multi-stage patch matching by iteratively segmenting images into smaller patches and adjusting their scales according to match results. While PATS mitigates limited input resolution, it still involves unnecessary feature comparison, which restricts matching accuracy.
Following the coarse-to-fine matching idea, we emphasize the need to establish an improved intermediate search space to address the accuracy challenges in current feature matching methods. When designing a matching search space, two essential aspects warrant consideration: the ease of constructing the search space reliably and the precision of following matching within it. We propose the incorporation of semantics into the search space design to effectively tackle these dual concerns. Previous work [16, 17] have introduced semantic into feature matching to leverage the semantic invariance against matching noises. Nonetheless, they maintain the original search space and utilize semantic to enhance patch or keypoint features, leading to conflicting fine-grained search space and semantic labeling. In other words, current semantic perception methods struggle to accurately define boundaries between different semantics [18]. However, these boundaries are precisely where matching search spaces tend to cluster, due to significant changes in image features. Thus, achieving fine-grained search spaces enhanced by semantics can be prone to semantic errors. Conversely, we suggest a coarse yet robust search space, termed as semantic area matches (Fig. 1), to enable better integration with semantics.
This area-level search space determines its size according to internal semantic cues, ensuring the inclusion of an adequate amount of semantic information within a condensed area to differentiate it from the entire image. Specifically, we propose two typical semantic areas: one encompassing an entire object and the other representing the intersection of multiple semantic entities. Thus, benefiting from semantic robustness in matching, the search space can be established without being influenced by semantic errors at the boundaries. Meanwhile, irrelevant feature computation is circumvented in the matched semantic areas, thus allowing for high resolution input and improving the accuracy of following matching phases. To actualize this search space, we introduce Semantic Area Matching (SAM) to detect and match semantic areas across images, powered by the advanced Large Language Model-based (LLM) semantic segmentation method [18]. With precise semantic segmentation of images, we show that semantic area matches can be easily achieved by hand-crafted semantic features, leading to a notable enhancement in matching accuracy.
Due to the image-level size of semantic area matches, they can serve as direct inputs for point matching, similar to overlap area match [14]. However, semantic area matches are more refined than the overlap area. It converts the original point matching task into multiple point matching tasks. We propose a matching framework, named Area to Point Matching (A2PM, Fig. 2 top), to formulate the matching process with semantic area matches. Initially, It establishes semantic area matches between images and then extracts these areas from the original images to perform point matching within them. This framework offers several advantages: 1) By re-cropping matched areas from high-resolution images, point matching benefits from more detailed inputs than the original. 2) The decoupling of the search space establishment from point matching allows that the accuracy of various point matching methods can be improved by the same area matching method.
@@
However, there is no free launch in feature matching when leveraging semantic. Although SAM effectively detects and matches most semantic areas between images, the inherent abstraction property in semantics overlooks local details. This can lead to semantic ambiguity during matching, particularly when distinct instances coexist within the image. Therefore, SAM may identify doubtful areas that cannot be confidently matched. Besides, the semantic ambiguity may also lead to erroneous area matches in SAM, adversely affecting feature matching. To tackle this challenge, we turn to the intrinsic geometric properties of area matches. In particular, considering that area matches across images signifies the same 3D entity, the constraint of epipolar geometry naturally applies. Moreover, area matches in the same image pair must adhere to the same constraint, i.e. geometric consistency, which can be employed to address the semantic ambiguity. Hence, we establish the geometric consistency of area matches by utilizing the epipolar geometry constraint of point matches within these areas. It enables the proposed Geometry Area Matching (GAM) to integrate a point matcher for accurate refinement of area matches. In practice, the GAM first predicts true area matches from doubtful candidates generated by SAM (GAM Predictor, GP). Subsequently, all area matches undergo filtering by the GAM Rejector (GR) to identify superior matches with geometric consistency. The point matches within the accurate area matches are obtained at the same time. Furthermore, to handle the correspondence aggregation issue in less-semantic scenes, the Global Match Collection module (GMC) is incorporated in GAM, which involves collecting additional point matches globally based on the geometry consistency of inside-area correspondences. The GMC ensures the generation of well-distributed matches, advantageous for downstream tasks. Through the combination of SAM and GAM, our Semantic and Geometry Area Matching (SGAM, Fig. 2 bottom) is capable of achieving accurate area and point matching between images.
In sum, our contributions are as follows:
@@
    Introduction of a semantic-friendly intermediate search space for feature matching, called semantic area matches, accompanied by a corresponding matching framework named A2PM. This framework involves the initial establishment of semantic area matches between images, and then performs point matching within these area image pairs, improving the matching accuracy ultimately.
@@
    Utilizing LLM-based semantic segmentation method, SGAM shows consistent improvement on matching accuracy for sparse, semi-dense and dense matching methods, thereby leading to impressive pose estimation performance on various indoor and outdoor datasets.
@@ II Related Work
II-A Sparse Matching
Sparse matching relies on detected keypoints and their descriptors [5, 19]. Through the nearest neighbor search based on descriptor distances, point matches can be established between images. In the age of deep learning, recent work [8, 7, 20, 21] utilize deep CNN to achieve better learning feature. Specifically, SuperPoint [7] is early in providing feature detection and description networks and outperforms conventional methods. Subsequent work [22, 8, 6] leverage a unified network to detect and describe feature. At the same time, detached learning detection [23, 24, 25] and description [26, 27, 28] are proposed as well. After feature detection, point match searching and outlier rejection are also advanced by recent learning methods [29, 30, 31, 32, 33, 34]. Essentially, this framework relies on extremely fine-grained search space establishment, i.e. the feature points, to achieve accurate point matching. However, feature point detection poses significant challenges in scenes with low texture, repetitive patterns, extreme changes in illumination and scale, which ultimately leads to a decline in performance. In contrast, our method appropriately reduces the search space to semantic area matches, which are robust due to their semantic invariance.
II-B Semi-dense and Dense Matching
In order to avoid detection failure, semi-dense and dense framework is proposed [11, 9, 35], aiming at jointly trainable feature detection, matching and outlier rejection to establish point matches directly from image pairs. Initially, 4D CNN is used to extract and compare image features densely [11, 36, 37] in dense methods. Recent DKM [12] constructs a Gaussian Process using CNN and achieves leading performance. Owing to limited receptive field [35] of CNN, COTR [35] incorporates Transformer [38] to process dense feature extracted by CNN. To alleviate the high computational cost, semi-dense methods such as LoFTR [9] and its variants [39, 10] suggest selecting sparse patch matches after dense computation to refine point matches. Nonetheless, the dense feature comparison in these methods leads to redundant computation, thus limiting the resolution of input images. Furthermore, the redundant search space introduces noise from non-overlapping areas in the image pair, resulting in degraded overall precision. As a solution, we propose to utilize semantics to reduce redundant computation through our semantic area matches. Due to the higher resolution and less noise in the area matches than original images, the precision of subsequent inside-area point matching is improved.
@@ II-C Coarse-to-Fine Matching
Recent matching methods commonly adopt a coarse-to-fine matching approach, which is deemed as a consensus in the field. Within this hierarchical process, intermediate search spaces play a critical role in determining the overall matching accuracy, and these spaces vary across different methods. Sparse methods typically identify a keypoint set as the search space for point matching, but struggle to achieve fine-grained keypoints. In semi-dense methods [9, 39, 10], patch-level matches are established through feature attention of the coarse level matching, which serve as the search space for the fine level matching. This approach significantly reduces time consumption compared to the dense counterpart [35]. On the other hand, dense methods [12, 36, 37] refine dense warps from coarse to fine feature maps, where dense correspondences between coarse feature maps can be viewed as dense patch matches between original images, making patch matches the intermediate search spaces for dense methods as well. Despite these advancements, the patch-level search spaces lack a clear association with image context, thus requiring expensive feature comparison in establishment. The accuracy issue in the coarse matching also persists due to low input resolution and error-prone redundant computations. While PATS [15] is proposed to extract more accurate features from equally cropped image patches with high resolution, the presence of redundant feature comparisons continues to limit matching accuracy. Recent overlap estimation methods [14, 13] utilize stage-one matching on the entire images to achieve the overlap between images. Then the stage-two point matching is performed inside the overlap area images. However, individual overlap area may be too coarse for precise point matching and the overlap establishment in these methods are expensive. Conversely, our method offers a semantic-aware search space, i.e. the semantic area matches. The incorporation of semantic enables robust establishment and efficient reduction of redundant computation of this search space, leading to impressive matching accuracy.
III Methodology
In this section, we first formulate the A2PM framework and its geometry properties (Sec. III-A). Then, we propose the Semantic Area Matching (Sec. III-B) with hand-crafted detection and description to establish putative semantic area matches, utilizing semantic robustness to overcome matching noises. To refine the area matches, we leverage geometry consistency formulated in Sec. III-A2 and propose the Geometry Area Matching (Sec. III-C). Finally, the implementation of A2PM framework by combining SGAM with any point matcher is illustrated (Sec. III-D). We provide a summarise of symbols used in Tab. VI-A of the appendix.
III-A Formulation
The formulation section includes the detailed description about the the proposed A2PM framework and the proposed geometry consistency of area matching.
III-A1 A2PM Framework
@@
Generally, given an image pair ($I_{0},I_{1}$), an area matching method $AM$ and a point matching method $PM$, the A2PM framework ($\mathcal{M}_{A}$) is responsible for connecting the area matching and point matching to achieve the final point matches accurately:
$\mathcal{P}=\mathcal{M}_{A}(I_{0},I_{1},AM,PM).$ (1)
@@
The output $\mathcal{P}=\{q^{m},p^{m}\}_{m}^{M}$ is the set of point matches ($q\in I_{0},~{}p\in I_{1}$ are correspondences). Specifically, the area matching can be formulated as follows. Suppose two matched areas in $I_{0},I_{1}$ are respectively $\{\alpha_{i}\}_{i}$ and $\{\beta_{j}\}_{j}$. The area match is represented as $\mathcal{A}_{i,j}=(\alpha_{i},\beta_{j})$. In area matching, $N$ pairs of area matches can be achieved by the $AM$:
$\{\mathcal{A}_{i,\pi(i)}\}_{i}^{N}=AM(I_{0},I_{1}),$ (2)
@@
where $\pi(i):\mathbb{R}\rightarrow\mathbb{R}$ is the index mapping between matched areas. Then, point matches inside each area match can be found to compose the final point matches:
$\mathcal{P}\leftarrow\{PM(\mathcal{A}_{i,\pi(i)})\}_{i}^{N}.$ (3)
@@
Due to the higher resolution and less redundancy of semantic area matches than the original input images of $PM$, the accuracy of final point matches is improved.
III-A2 Geometry Consistency of Area Matching